# Complex Reasoning over Logical Queries on Commonsense Knowledge Graphs

### v1

@@ 7 Conclusion
In this paper, we leverage the concept of conjunctive logical queries to create a complex commonsense reasoning dataset derived from CSKGs. The dataset, Com, comprises a human-annotated evaluation set and a distantly supervised training set without further annotations. Our experiments highlight the challenging nature of complex commonsense reasoning that involves multiple events or multi-hop scenarios, even for advanced language models such as GPT-4. Additionally, we train question answering models and generative commonsense reasoning models using Com. The results show significant improvements across eight diverse downstream commonsense reasoning tasks, highlighting the potential of leveraging CSKGs to acquire complex reasoning signals inexpensively, without relying on extra human effort.
@@
The construction of Com relies on sampling complex logical queries from existing CSKGs, which requires addressing sparsity, quality, contextualization issues. Despite conducting normalization and filtering, there may still be missing links within ATOMIC and mislabeled or ambiguous triples, which limits the quality of our sampled queries. Future works can focus on deriving complex queries from CSKGs with better quality and more diverse semantics, which should also have higher density, such as on ATOMIC-10x, NovATOMIC West et al. (2023).
@@
We sample the data from ATOMIC, which is an open-source commonsense knowledge graph that may contain biases around gender, occupation, and nationality Mehrabi et al. (2021). When constructing Com, these biases may propagate if biased triples are sampled in a complex query that becomes of the training set. We collected 1.3k inferences through crowdsourcing. The participants were compensated with an hourly wage of 16 USD, which is comparable to the minimum wages in the US. The qualification was purely based on the workers’ performance on the evaluation set, and we did not collect any personal information about the participants from MTurk.
@@ A.1 Nodes Normalization (Dealing with Sparsity)
To alleviate the sparsity issue, we first normalize the tail entities. In ATOMIC, heads are pre-defined complete sentences (for example, “PersonX says sorry”) while tails are usually short phrases without a subject (for example, “to say sorry”). This discrepancy produces many duplicated nodes and make the graph sparser. We develop simple rules to add “PersonX” or “PersonY” in front of the tails to make them a complete sentence, if the tail does not have a subject. This process merged 3.7% nodes together. We present the normalization rules in Table 6.
Second, as the nodes in ATOMIC are free-text, some nodes with the same semantic meaning are represented as separated nodes due to some minor annotation distinctions and errors, e.g., “PersonX buys a ticket” versus “PersonX buys a ticket .”. We use a state-of-the-art sentence embedding model444https://huggingface.co/sentence-transformers/all-mpnet-base-v2, to merge nodes with cosine similarity score over 0.95. In this process, 20.0% nodes are merged together and the average degree increases by 25.3%.
| Relations |
Mapping rules |
Add PersonX/Y in front of the tail and remove the initial “to” |
xEffect/oEffect |
Add PersonX/Y in front of the tail |
xReact/oReact |
Add PersonX/Y and “is” in front of the tail |
xAttr |
Add a PersonX/Y and “is” in front of the tail |
@@ A.3 Verbalization
Query Type |
Question Template |
2i |
What event or state is both Prompt(r1) [V1] and also prompt(r2) [V2]? |
3i |
What event or state is both Prompt(r1) [V1], Prompt(r2) [V2], and also Prompt(r2) [V2]? |
2p |
What event or state is Prompt(r1) {Prompt(r2) [V1]}? |
ip |
What event or state is prompt(r3) {both prompt(r1) [V1], and also prompt(r2) [V2] }? |
pi |
What event or state is both prompt(r1) {prompt(r3) [V3]}, and also prompt(r2) [V2]? |
Relation |
Prompt Template |
xIntent |
the intention of PersonX before |
xNeed |
what PersonX needed to do before |
xWant |
what PersonX wants to do after |
xEffect |
the effect on PersonX after |
xReact |
what PersonX feels after |
xAttr |
what PersonX is seen as given |
oEffect |
the effect on PersonY after |
oReact |
what PersonY feels after |
oWant |
what PersonY wants to do after |
HinderedBy |
what hindered |
isAfter |
what happens before |
isBefore |
what happens after |
@@ A.4 Human Annotation
Vera |
[Context] [Question] [Option] |
HyKAS, CAR |
[Context] [Question] [Option] |
Model |
Prompt |
@@ Appendix C Additional Analysis
In contrast, our complex commonsense reasoning benchmark introduces additional complexities compared to ParaCOMET and COMET-M. Besides the complex structures in the context that involves multiple events, the desired relation or question involves multi-hop reasoning as well. For instance, rather than focusing on the cause of a single sentence or event, Com explores questions related to common causes, effects, attributions of multiple events, and two-hop inferences. This distinctive formulation sets our work apart and poses a greater challenge for LLMs to effectively reason and provide accurate responses.
@@ C.1 Difficulty of Different Query Types
| Type |
Context |
Question |
COMET |
Com-COMET |
| 2p |
Ezra updates Ezra’s resume (V1) |
What event or state is the intention of Ezra before the intention of Ezra before V1? |
be financially independent ✓ |
Every day, Benjamin goes to work diligently (V1), never missing a day. They are dedicated and committed to their job. In particular, Benjamin doesn’t work hard on it (V2) and instead takes a more relaxed approach, focusing on maintaining a healthy work-life balance. |
What event or state is both the effect on Benjamin after Benjamin go to work every day (V1) and also what hindered Benjamin work hard on it (V2)? |
Benjamin is sick ? (Not perfect as Benjamin is trying to keep a work-life balance instead of having a sick leave) |
Benjamin gets tired from working hard ✓ |
| 2i |
Chloe is known for being hardworking (V1) and dedicated. As a result, Chloe leads a good life (V2). |
What event or state is both the effect on Chloe after Chloe is hardworking (V1) and also what Chloe wants to do after Chloe leads a good life (V2)? |
to have a good life ? (No inferential gap) |
to have success in life ? (No inferential gap) |
| ip |
After looking for a new car (V1), Lydia is driving to school (V2). |
What event or state is what Lydia needed to do before the event that is both what Lydia wants to do after Lydia is looking for a new car (V1), and also what Lydia needed to do before Lydia is driving to school (V2)? |
None ✗ |
take a car for test drive ✓ |

### v2

@@ 7 Conclusion
In this paper, we leverage the concept of conjunctive logical queries to create a complex commonsense reasoning dataset derived from CSKGs. The dataset, Com2, comprises a human-annotated evaluation set and a distantly supervised training set without further annotations. Our experiments highlight the challenging nature of complex commonsense reasoning that involves multiple events or multi-hop scenarios, even for advanced language models such as GPT-4. Additionally, we train question answering models and generative commonsense reasoning models using Com2. The results show significant improvements across eight diverse downstream commonsense reasoning tasks, highlighting the potential of leveraging CSKGs to acquire complex reasoning signals inexpensively, without relying on extra human effort.
Acknowledgement
Yangqiu Song was supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. Yangqiu Song thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08). We thank the support from the Tencent AI Lab Rhino-Bird Focused Research Program. We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Science Seed Fund, the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
@@ Acknowledgement
The construction of Com2 relies on sampling complex logical queries from existing CSKGs, which requires addressing sparsity, quality, contextualization issues. Despite conducting normalization and filtering, there may still be missing links within ATOMIC and mislabeled or ambiguous triples, which limits the quality of our sampled queries. Future works can focus on deriving complex queries from CSKGs with better quality and more diverse semantics, which should also have higher density, such as on ATOMIC-10x, NovATOMIC West et al. (2023).
@@
We sample the data from ATOMIC, which is an open-source commonsense knowledge graph that may contain biases around gender, occupation, and nationality Mehrabi et al. (2021). When constructing Com2, these biases may propagate if biased triples are sampled in a complex query that becomes of the training set. We collected 1.3k inferences through crowdsourcing. The participants were compensated with an hourly wage of 16 USD, which is comparable to the minimum wages in the US. The qualification was purely based on the workers’ performance on the evaluation set, and we did not collect any personal information about the participants from MTurk.
@@ A.1 Nodes Normalization (Dealing with Sparsity)
To alleviate the sparsity issue, we first normalize the tail entities with simple rules similar with that in Dense-ATOMIC Shen et al. (2023) and CKBP Fang et al. (2021a). In ATOMIC, heads are pre-defined complete sentences (for example, “PersonX says sorry”) while tails are usually short phrases without a subject (for example, “to say sorry”). This discrepancy produces many duplicated nodes and make the graph sparser. We develop simple rules to add “PersonX” or “PersonY” in front of the tails to make them a complete sentence, if the tail does not have a subject. This process merged 3.7% nodes together.
Second, as the nodes in ATOMIC are free-text, some nodes with the same semantic meaning are represented as separated nodes due to some minor annotation distinctions and errors, e.g., “PersonX buys a ticket” versus “PersonX buys a ticket .”. These discrepencies can be addressed using embedding similarities Wu et al. (2023). We use a state-of-the-art sentence embedding model444https://huggingface.co/sentence-transformers/all-mpnet-base-v2, to merge nodes with cosine similarity score over 0.95. In this process, 20.0% nodes are merged together and the average degree increases by 25.3%.
| Relations | Mapping rules | ||
Add PersonX/Y in front of the tail and remove the initial “to” | ||
| xEffect/oEffect | Add PersonX/Y in front of the tail | ||
| xReact/oReact | Add PersonX/Y and “is” in front of the tail | ||
| xAttr | Add a PersonX/Y and “is” in front of the tail |
@@ A.3 Verbalization
| Query Type | Question Template |
| 2i | What event or state is both Prompt(r1) [V1] and also prompt(r2) [V2]? |
| 3i | What event or state is both Prompt(r1) [V1], Prompt(r2) [V2], and also Prompt(r2) [V2]? |
| 2p | What event or state is Prompt(r1) {Prompt(r2) [V1]}? |
| ip | What event or state is prompt(r3) {both prompt(r1) [V1], and also prompt(r2) [V2] }? |
| pi | What event or state is both prompt(r1) {prompt(r3) [V3]}, and also prompt(r2) [V2]? |
| Relation | Prompt Template |
| xIntent | the intention of PersonX before |
| xNeed | what PersonX needed to do before |
| xWant | what PersonX wants to do after |
| xEffect | the effect on PersonX after |
| xReact | what PersonX feels after |
| xAttr | what PersonX is seen as given |
| oEffect | the effect on PersonY after |
| oReact | what PersonY feels after |
| oWant | what PersonY wants to do after |
| HinderedBy | what hindered |
| isAfter | what happens before |
| isBefore | what happens after |
@@ A.4 Human Annotation
| Vera | [Context] [Question] [Option] | |||||
| HyKAS, CAR | [Context] [Question] [Option] |
| Model | Prompt | |||
@@ Appendix C Additional Analysis
In contrast, our complex commonsense reasoning benchmark introduces additional complexities compared to ParaCOMET and COMET-M. Besides the complex structures in the context that involves multiple events, the desired relation or question involves multi-hop reasoning as well. For instance, rather than focusing on the cause of a single sentence or event, Com2 explores questions related to common causes, effects, attributions of multiple events, and two-hop inferences. This distinctive formulation sets our work apart and poses a greater challenge for LLMs to effectively reason and provide accurate responses.
@@ C.1 Difficulty of Different Query Types
| Type | Context | Question | COMET | Com2-COMET | ||
| 2p | Ezra updates Ezra’s resume (V1) | What event or state is the intention of Ezra before the intention of Ezra before V1? |
be financially independent ✓ | ||
Every day, Benjamin goes to work diligently (V1), never missing a day. They are dedicated and committed to their job. In particular, Benjamin doesn’t work hard on it (V2) and instead takes a more relaxed approach, focusing on maintaining a healthy work-life balance. | What event or state is both the effect on Benjamin after Benjamin go to work every day (V1) and also what hindered Benjamin work hard on it (V2)? | Benjamin is sick ? (Not perfect as Benjamin is trying to keep a work-life balance instead of having a sick leave) | Benjamin gets tired from working hard ✓ | ||
| 2i | Chloe is known for being hardworking (V1) and dedicated. As a result, Chloe leads a good life (V2). | What event or state is both the effect on Chloe after Chloe is hardworking (V1) and also what Chloe wants to do after Chloe leads a good life (V2)? | to have a good life ? (No inferential gap) | to have success in life ? (No inferential gap) | ||
| ip | After looking for a new car (V1), Lydia is driving to school (V2). | What event or state is what Lydia needed to do before the event that is both what Lydia wants to do after Lydia is looking for a new car (V1), and also what Lydia needed to do before Lydia is driving to school (V2)? | None ✗ | take a car for test drive ✓ |